OpenStudy (anonymous):

-5x+y=-5 5x-y=5 Does this system have a solution?

OpenStudy (nikato):

no?

OpenStudy (anonymous):

No, both variables cancel out

OpenStudy (anonymous):

what about this: 3x−y=3ˏ −3x+y=3

OpenStudy (anonymous):

Isolate y for both instances. We get 5x-5=y. This means the two lines lie on top of each other. That should mean that x is all reals.

OpenStudy (anonymous):

Same thing

OpenStudy (anonymous):

Once again you must isolate for y. You get 3x-3=y and 3x+3. Since the slope is the same the two lines are parallel and cannot intersect at any point. No solution.

OpenStudy (nikato):

oh, nvm, the first one is infinitely many

OpenStudy (anonymous):

Yeah i just realized that